A stone is dropped from top of a tower. When it crosses a point 5m below the top, another stone is let to fall from a point 25m below the top. Both stones reach the bottom of tower simultaneously.Find the height of the tower.(g=10 m/s2)

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Chilukuri

tami taken to fall a distance 5 m by first stone

5=1/2 *g*t2

t=1 sec

let hight of tower is h

so motion of first stone

 h= 1/2 * 10 * t2              .....................1

motion of second stone

h-25 =1/2 * 10 *(t-1)2                   ...............2

put value of h from equation 1

 t=3 sec

and h=45 m


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin